Two refills for a ball point pen are selected at random for a box that contains 55 blue, 3 red and 3 green refills. If X is the number of blue refills and Y is the number of red refills selected .Find the joint probability function f (x, y) and p[(x,y)∈A ], where A is the region {(x,y): x+y<=1}


1
Expert's answer
2021-08-17T14:26:07-0400

Part 1

The possible pairs of values (x, y) are (0, 0), (0, 1), (1, 0), (1, 1), (0, 2), and (2, 0), where p(0, 1), for

example represents the probability that a red and a green refill are selected. The total number of equally likely ways of selecting any 2 refills from the 8 is

"\begin{pmatrix}
 61 \\
 2 
\end{pmatrix}= \frac{61!}{2!58!}=107970"

The number of ways of selecting 1 red from 3 red refills and 1 green from 3 green refills is

"\begin{pmatrix}
 3 \\
 1
\end{pmatrix}\begin{pmatrix}
 3 \\
 1
\end{pmatrix}=9"

Hence, "p(1, 1) = \frac{9}{107970}" . Similar calculations yield the probabilities for the other cases, which are presented in the following table. Note that the probabilities sum to 1.

"p(x,y)= \frac{\begin{pmatrix}
 55 \\
 x
\end{pmatrix}\begin{pmatrix}
 3 \\
 y
\end{pmatrix}\begin{pmatrix}
 3 \\
 3-x-y
\end{pmatrix}}{\begin{pmatrix}
 61 \\
 2
\end{pmatrix}}\\
For \space x =0,1,2\\
For \space y =0,1,2\\
0≤x+y≤2"


Part 2

"P[(X, Y) \in A] = P(X + Y≤1)\\
= p(0, 0) + p(0, 1) + p(1, 0)\\
=\frac{55}{107970}+\frac{3}{53985}+\frac{165}{107970}\\
=\frac{113}{53985}"